The length of a rectangular lot is 50 feet more than the width. If the perimeter is 500 feet, then what are the length and width?
How do I solve this?

Answer by ankor@dixie-net.com(15645)

Write an equation for each statement/phrase:
:
"length of a rectangular lot is 50 feet more than the width."
L = W + 50
:
I assume you know the perimeter equation:
"If the perimeter is 500 feet,"
2L + 2W = 500
:
then what are the length and width?
:
If we could get the above equation to be a single unknown it would be easy to
solve. The 1st equation says that L = (W+50), substitute (W+50) for L in the
perimeter equation:
2(W+50) + 2W = 500
:
We can simplify this equation by dividing thru by 2, then we have:
(W+50) + W = 250; a very simple equation to solve for W
:
2W = 250 - 50
2W = 200
W = 200/2
W = 100 ft is the width;
:
Remember: "length of a rectangular lot is 50 feet more than the width."
L = 100 + 50
L = 150
:
Check our solutions in the perimeter equation:
2(150) + 2(100) = 500
;
How about this? Comprendez?